OpenStudy (anonymous):

what multiplies to -64 and add to 64

OpenStudy (anonymous):

Please stop posting the same question over and over.

Directrix (directrix):

(-64)/x + x = 64 x = 8 (4 + √17) and 8( 4 - √17) See if those two work.

Directrix (directrix):

xy = -64 so y = (-64)/x and y + x = 64